OpenStudy (anonymous):

Simplify: 8 – 5 + 3 • 6

OpenStudy (anonymous):

Seriously? Thats it?

OpenStudy (anonymous):

the dot means 'multiply' so it should be: 8-5+(3*6) =21 correct me if i'm wrong

OpenStudy (anonymous):

8-5+18 Use BODMAS

OpenStudy (maheshmeghwal9):

YA AS Harkirat.:- 8-5+18 =>8+13 =>21.

OpenStudy (maheshmeghwal9):

ORDER:------- Division, Multiplication, Addition, Subtraction. K :) --------->BODMAS Rule.

OpenStudy (anonymous):

thank you so much:)

OpenStudy (maheshmeghwal9):

yw:)

OpenStudy (anonymous):

how did you get 8 +13 though

OpenStudy (maheshmeghwal9):

Put 8 one side. Now how I get 13? The answer is that -5+18=18-5=13.

OpenStudy (maheshmeghwal9):

so it becomes 8+13. As u see.

OpenStudy (maheshmeghwal9):

in BODMAS rule, it is given that u have to add first then subtract But it is ur choice What to do first:) u can do anything first but in only choice of add & subtract:)

OpenStudy (maheshmeghwal9):

tell me where is ur doubt?

OpenStudy (maheshmeghwal9):

I can help u more:)

OpenStudy (anonymous):

ok i get now:)

OpenStudy (maheshmeghwal9):

are u a newcomer to open-study?

OpenStudy (maheshmeghwal9):

so I tell u one thing that u have to close this question firstly if u have got it; to ask a new one:)

OpenStudy (maheshmeghwal9):

& by the way in which class do u study?

OpenStudy (anonymous):

yes im new here lol

OpenStudy (anonymous):

and its algebra

OpenStudy (maheshmeghwal9):

k!

OpenStudy (maheshmeghwal9):

best of luck for ur studies:)